S   +   6 HNO3 →  H2SO4  +  6 NO2  +   2 H2O

In the above equation, how many grams of water can be made when 12.6 moles of HNO3 are consumed?

Round your answer to the nearest tenth. If you answer is a whole number like 4, report the answer as 4.0

Use the following molar masses. If you do not use these masses, the computer will mark your answer incorrect.:


1
Expert's answer
2022-02-06T11:20:04-0500

The molar mass of H2O is 18.0 g/mol.


"m(H_2O)=12.6\ mol(HNO_3)\times\frac{2\ mol(H_2O)}{6\ mol(HNO_3)}\times\frac{18.0\ g(H_2O)}{1\ mol(H_2O)}=75.6\ g"


Answer: 75.6 g